OpenStudy (anonymous):

Amy has a garden covering 900 square feet. a. If she wants to plant moss roses on 10% of the garden, how many square feet will moss roses cover? Write and solve a percent equation to solve this problem. b. If the garden occupies 25% of Amy’s yard, what is the area of her yard in square feet? Write and solve a percent equation to solve this problem.

OpenStudy (anonymous):

GIVEN: Amy has a garden covering 900 square feet. a. If she wants to plant moss roses on 10% of the garden, how many square feet will moss roses cover? Write and solve a percent equation to solve this problem. b. If the garden occupies 25% of Amy’s yard, what is the area of her yard in square feet? Write and solve a percent equation to solve this problem.

OpenStudy (anonymous):

FIND:?????

OpenStudy (anonymous):

G=900SQ.FT.

OpenStudy (anonymous):

900SQ.FT. - 10% OF G=SQ.FT. OF MOSS IN G.....

OpenStudy (anonymous):

HENCE X=10% OF 900

OpenStudy (anonymous):

9.....

OpenStudy (anonymous):

HENCE 900-9=ANS.....

OpenStudy (anonymous):

Ans.....891!

OpenStudy (anonymous):

do the rest

OpenStudy (anonymous):

woah @_@ umm okay im confused

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got this for A. Equation: x/900 = 1/100 x = 90 sq. ft. is that correct?

OpenStudy (anonymous):

read my steps

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, what about b.? I don't know what to do for that

OpenStudy (anonymous):

x/900=2.5/100

OpenStudy (anonymous):

okay, thanks